OpenStudy (anonymous):

if A is a nilpotent matrix of order p,then find the inverse of the matrix I+A and I-A

OpenStudy (anonymous):

@UnkleRhaukus come help this brother here; I forgot how to do matrix

OpenStudy (unklerhaukus):

i can't remember linear algebra at the momnet

OpenStudy (anonymous):

Me neither.

OpenStudy (loser66):

According to this site, near bottom, \[(I+A)^{-}= I-N+N^2-N^3+....\] http://en.wikipedia.org/wiki/Nilpotent_matrix

OpenStudy (loser66):

and according to my note, my prof said that \[(I-N)^{-}=I+N+N^2+N^3+....\]

OpenStudy (loser66):

his argument is \[\frac{1}{x}=1+x+x^2+x^3+....\] \[\frac{1}{I-A}=(I-A)^{-} = I+A+A^2+A^3+...\]like what I write above. hehehe...everything doesn't relate to me.

OpenStudy (anonymous):

\[\left[ I-A \right]^{-1}=\frac{ I }{ I-A }=I+A+A^2+...+A^{p-1}+\frac{ A^p }{ I-A }\]and \[A^p=[0]\]Then,\[\left[ I-A \right]^{-1}=I+\sum_{k=1}^{p-1}A^k\]